a^2−8ab+16b^2=81, find a–4b
Answer by fractalier(6550)

We can factor
a^2 - 8ab + 16b^2 = 81
and get
(a - 4b)^2 = 81
Now taking the square root of both sides we get
a - 4b = 9 or
a - 4b = -9
Two solutions.